OpenStudy (anonymous):

1 2 4 3 6 8 7 14 16? whats the answer

OpenStudy (anonymous):

whats the question?????

OpenStudy (anonymous):

whats the next sequence ?

OpenStudy (anonymous):

the number that best completes the sequence of 1 2 4 3 6 8 7 14 16 ?

OpenStudy (anonymous):

1*2=2 2+2=4 4-1=3 3*2=6 6+2=8 8-1=7 7*2=14 14+2=16 16-1=15 15*2=30 30+2=32 32-1=31

OpenStudy (anonymous):

so the answer is 31 or 32???

OpenStudy (anonymous):

nope its 14

OpenStudy (anonymous):

sorry 15

OpenStudy (anonymous):

1 2 4 3 6 8 7 14 16,15,

OpenStudy (anonymous):

multiply by 2 add 2 subtract 1 thats the rule

OpenStudy (anonymous):

thanks i need help on this one also 10 30 15 16 48 24 25?

OpenStudy (anonymous):

2+2=4 4-1=3 3*2=6 6+2=8 8-1=7 7*2=14 14+2=16 16-1=15

OpenStudy (anonymous):

so i just do all that?? cus i just need to really find the answer??

OpenStudy (anonymous):

triple and half

OpenStudy (anonymous):

10 30 15 16 48 24 25 triple and half

OpenStudy (anonymous):

10 30 15 16 48 24 25? multiply by 3 divide by 2 add 1

OpenStudy (anonymous):

thats the rule

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

ok so i did 25*3=75 then i divided it by 2 but i got 37R1

OpenStudy (anonymous):

ok 75 is correct 75/2=?

OpenStudy (anonymous):

10*3=30(multiply by 3) 30/2=15(divide by 2) 15+1=16(add 1) 16*3=48(multiply by 3) 48/2=24(divide by 2) 24+1=25(add 1) 25*3=75(multiply by 3)

OpenStudy (anonymous):

2 9 5 13 10 19 17?

OpenStudy (anonymous):

i need the answer to this one as well

OpenStudy (anonymous):

Ok,guys, cool it.....

OpenStudy (anonymous):

Keep it civil.

OpenStudy (anonymous):

I mean it....

OpenStudy (anonymous):

ok i will be the more wiser one and stop

OpenStudy (anonymous):

Smart move...

OpenStudy (anonymous):

Sorry for all that nonsense. Please feel free to continue your question.

OpenStudy (anonymous):

ok thank you for the help tho

OpenStudy (anonymous):

ok sequences are hard for me so i need help with this one also 3 9 4 5 25 20 21 441?

OpenStudy (anonymous):

need help finding the answer??